OpenStudy (kingswank):

MEDALS WILL BE GIVEN!!! Use the net to find the surface area of the rectangular prism. With the measurements of 6ft,11ft,6ft,11ft & height of 8ft.

OpenStudy (anonymous):

Find the volume and surface area of the given solids: 1). Cube - l = 2 in, w =2 in, and h = 2 in V = 8 cubic inches SA = 24 square inches 2). Cube - l =1 in, w = 1 in, and h = 1 in V = 1 cubic inch SA = 6 square inches 3). Prism - l = 7 in, w = 4 in, and h = 11 in V = 308 in^3 Surface Area of Prism: [ 2(lw) + (2l + 2w)h ] = 56 + 22*11 = 298 square inches 4). Rectangular prism - l = 5 in, w = 4 in, and h = 3 in V = 60 in^3 SA (Cuboid): [ 2(lw) + (2l + 2w)h ] = 40 + 18*3 = 94 square inches 5). Pyramid Volume =1/3 x area of base x height Surface Area of Any Pyramid = area of base + area of each of the lateral faces Regular pyramid = area of base + 1/2 ps where p = perimeter of the base, s = slant height Square pyramid = b^2 + 2bs where b = length of the base, s = slant height Square pyramid - l = 8 in, w = 8 in, and h = 10 in s = sqrt (164) V = 213 1/3 in^3 SA = 64 + 16*(164)^(1/2) = 268.89997559785115 in^2

OpenStudy (kingswank):

Is that a question or are you helping me?